OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence? A. 14/13 B. 25/7 C. 4

OpenStudy (mathstudent55):

To find term \(a_n\), use this formula: \(a_n = a_1 + (n - 1)d\) You have \(a_1 = -3\), \(n = 15\), \(a_{15} = 53\)

OpenStudy (mathstudent55):

Substitute those values in the formula, and solve for d.

OpenStudy (anonymous):

x = a + d(n-1) 53 = -3 +d(15-1) 56 = 14d